# The Jungle of Generative Drug Discovery: Traps, Treasures, and Ways Out

### v1

@@ Abstract
Temperature has the greatest impact on the characteristics of de novo designs (Fig. 5). Higher temperature values increase the design diversity across datasets (Fig. 5a,c), but reduce validity and similarity to the fine-tuning set (Fig. 5a,b), in agreement with previous works9, 26. For top-$k$ sampling, The smallest value of k ($k=3$) causes mode collapse, i.e., designs are valid but repetitive (Fig. 5a), suggesting that considering three candidate tokens might be insufficient to design diverse molecules. For $k\geq 5$, the sampling behavior approximates temperature sampling with $T=1$, indicating that the top-5 tokens already cover the most likely tokens. Top-$p$ sampling behaves similarly to top-k sampling. For $p\leq 0.9$, designs are valid but repetitive, while larger values of $p$ resemble temperature sampling ($T=1$). In the former case, few tokens cover the $p$ threshold, causing mode collapse; in the latter case, adding more tokens has little effect due to their low probability. We call this behavior, which is consistent across model architectures (Fig. S7, S8), ‘filtering trap’.
@@
These findings contrast with natural language generation: while top-$k$ and top-$p$ outperform temperature sampling to generate natural language, temperature sampling is the primary sampling strategy to control molecule diversity. To gain further insights, we generated 102,400 designs (using the fine-tuned LSTM for VDR) and analyzed the frequency of appearance of each token in the top-5 and top-90% of the distribution (Fig. 6a,b), along with their mean sampling probability (Fig. 6c). While probable tokens, such as carbon, are consistently likely across generation steps, most tokens rarely appear among the top-5 and top-90%. This skewed distribution differentiates molecule design from natural language generation, where tokens can be sampled among hundreds of thousands of candidates. In the molecular field, the number of tokens is inherently constrained due to the number of elements that can be used in drug-like molecules, leading to a higher concentration of likely candidates and a less diverse set of options at each generation step. Overall, our analysis reveals a unique behaviour of molecule generation, and exemplifies the gaps between natural language and drug discovery applications.
@@
Table 1: Traps, treasures, and ways out. Summary of identified the pitfalls, robust tools, and guidelines proposed in this study. These considerations were divided based on the evaluation stage they pertain to.
Stage Trap Treasure Way Out
@@
Library similarity Library size. Metrics like FCD and FDD are dependent on library size, and decrease with increasing number of designs. FCD. FCD plateaus for more than 10,000 designs, making it suitable even with few reference bioactive molecules. Evaluating large libraries. Report FCD values for more than 10,000 designs, and FDD values for more than 1000 designs.
@@
Molecule generation Token filtering. Considering a small token subset during molecule sampling (via top-k or top-p sampling) cause mode collapse (repetitive, low-diversity designs). Temperature sampling. Temperature sampling is the most effective strategy to control diversity and balance novelty vs. validity. Varying T values. Controlling diversity through token subsets is ineffective due to the unique behavior of molecule sampling. Temperature sampling should be used, with varying $T$ to tune lead optimization vs. diversity-focused exploration.
Orienteering in the jungle of generative drug discovery
Generative deep learning is reshaping computer-assisted drug discovery, and we have only begun to test its potential and boundaries. In this context, robust evaluation pipelines are essential to identify and eventually expand these boundaries. Thanks to our large-scale analysis, our work highlights previously overlooked challenges and untapped opportunities in the application of generative deep learning to drug discovery. These findings, in turn, highlight new solutions for evaluating generative drug discovery models (Table 1).
One of the most important findings of this study is that the number of the generated designs can act as a misleading confounder in quality analysis – an issue we term the ‘size trap’. Such trap has significant implications: similarity and diversity metrics must be compared across libraries of consistent sizes. We also advocate for reporting these metrics for more than 105 designs. To adddress the computational challenges of diversity assessment of large libraries, we introduce the number of substructures as a fast and reliable metric.
This study also bridges knowledge gaps and highlights barriers in the quest to ‘tame the jungle’ of generative drug discovery. By strengthening the connections between drug discovery and natural language processing 60, we emphazise the usefulness of design likelihoods to select molecules for prospective studies, remove “recurring hallucinations”, and rationalize the exploration-exploitation trade-off. Additionally, we uncover key differences between the two fields in the generation phase, revealing characteristic properties of molecule generation. In this context, we recommend temperature sampling as an effective strategy to navigate the chemical space with generative deep learning approaches.
Drug discovery is inherently a multi-objective endeavor, making the definition and measurement of design quality a persistent challenge. However, we expect that the insights and connections uncovered in this study will provide a solid foundation for making an informed progress towards pushing the current boundaries of generative drug discovery.
@@ Methods
Pre-training set. 2,372,675 SMILES strings were obtained from ChEMBL v33 44. Salts were removed and molecules composed only of C, H, O, N, S, P, F, Cl, Br, I atoms were retained. The SMILES strings of the remaining compounds were sanitized, canonicalized, and the charge and sterochemistry annotations were removed. SMILES strings longer than 80 tokens were dropped. The final set consisting of 1,584,858 molecules was randomly divided into training ($n$ = 1,500,000), validation ($n$= 40,000), and test splits ($n$ = 44,858).
Fine-tuning sets. The fine-tuning datasets were curated from ExCAPE-DB 45. The targets Dopamine Receptor D3 (DRD3), peptidyl-prolyl cis/trans isomerase (PIN1), and Vitamin D Receptor (VDR) were selected. The bioactive molecules available in the pret-training set were excluded, and the remaining molecules underwent the same pre-processing steps as described above. 320 training, 128 validation, and 128 test molecules were randomly sampled among the pre-processed strings. Random sampling was repeated five times, obtaining five fine-tuning sets per protein target.
@@
where $T$ is the temperature parameter and $y_{t}$ is the logit output by the model for the token $t$. Increasing the temperature value increases the uniformity of the distribution (uniform distribution for $T\rightarrow\infty$), while decreasing the temperature value decreases the randomness (Dirac distribution for $T\rightarrow 0$). When $T$ = 1, the so-called multinomial sampling (where the probability of generating each token is determined by the model output) is performed. We experimented with 0.5, 0.75. 1.0, 1.25, 1.5, 1.75, and 2.0 as $T$ values in this study.
Top-$k$ sampling. Top-$k$ sampling 57 samples the next token from the most likely $k$ tokens. Increasing the $k$ values to the number of tokens in the vocabulary makes it equivalent to temperature sampling. Using $k=1$ is equivalent to greedy sampling, i.e., sampling the most likely token at each step. We experimented with values of $k$ equal to 3, 5, 10, 15, 20, 25, and 30 (with a vocabulary size equal to 33).
Top-$p$ sampling. Top-$p$ sampling 57 (also known as nucleus sampling) samples the next token from the minimum cardinality set whose summed probabilities exceed the threshold $p$ ($0\leq p\leq 1$). Decreasing the value of $p$ includes fewer tokens in the selection and helps to avoid degenerate outputs, while trading diversity off 57. Top-$p$ sampling approximates greedy sampling as $p\rightarrow 0$ and is equivalent to temperature sampling when $p=1.0$. In this study, we experimented with values of $p$ equal to 0.5, 0.6, 0.7, 0.8, 0.9, 0.925, 0.950, and 0.975.
Sampling strategy. From each target (3x), each data split (5x), and each model architecture (3x), we generated 1,000,000 SMILES strings per sampling strategy and sampling parameter ($T$, $k$ and $p$, 22 values tested in total). This resulted in a total of $3\times 5\times 3\times 22\times 10^{6}=990,000,000=9.9\times 10^{8}\approx 10^{9}$ molecules that were designed and evaluated in this study.
@@
  • The Fréchet ChemNet Distance (FCD) 46 was computed using the fcd library released by the authors.
  • The Fréchet distance 48 between molecular descriptor distributions (FDD) was computed using the following descriptors: octanol-wated partitioning coefficient 61, molecular weight, number of hydrogen bond donors, number of rings, and topological surface area. Molecular descriptors were computed via rdkit, and min-max normalized to global maximum and minimum values before distance calculation.
  • The substructure similarity was computed via Tanimoto similarity on extended connectivity fingerprints (ECFPs) 50. ECFPs were computed with rdkit (fpSize=2048, and radius=2).
Internal diversity The number of clusters was computed by using the sphere exclusion algorithm implemented in the LeaderPicker module of rdkit. We used a distance threshold of 0.6, using the Tanimoto similarity on ECFPs. This is equivalent to computing #Circles metric 49. The number of substructures was calculated by counting the number of unique fingerprint keys identified by the Morgan algorithm (rdkit, radius=2).
@@
where $t$ runs over the tokens in the designed sequence. The log-sum-exp trick was used to mitigate numerical instabilities.
Data availability
The data used and generated in our study will be available on Zenodo upon paper acceptance and on GitHub at the following URL: https://github.com/molML/jungle-of-generative-drug-discovery.
Code availability
The Python code to replicate our study will be available on Zenodo upon paper acceptance and on GitHub at the following URL: https://github.com/molML/jungle-of-generative-drug-discovery.
Author Contributions
Conceptualization: all authors. Data curation: R.Ö. Formal analysis: all authors. Investigation: all authors. Methodology: all authors. Software: R.Ö. Visualization: R.Ö. Writing – original draft: R.Ö. Writing – review and editing: all authors.
Acknowledgements
This research was co-funded by the European Union (ERC, ReMINDER, 101077879). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The authors also acknowledge support from the Irene Curie Fellowship, the Centre for Living Technologies. The authors thank Selen Parlar and Helena Brinkmann for their feedback on the manuscript.
@@ Acknowledgements
The Jungle of Generative Drug Discovery: Traps, Treasures, and Ways Out
@@
Figure S1: Diversity metrics divided by library size. Number of substructures and clusters are computed in increasing library sizes, and divided by the number of molecules in the library. The same experimental and visualization setups are used as Figure 2.
Figure S2: Effect of synthesizability filtering on internal diversity. Synthetic accessibility of the designs are computed 62and the generations with a score above 6 are filtered out 62. Number of substructures is computed with the remaining compounds. The same experimental and visualization setups are used as Figure 2.
Figure S3: Speed comparison of internal diversity metrics. Number of clusters and substructures are computed in increasing library sizes, and number of seconds per computation is measured. Each computation is repeated 10 times and the average time passed is reported in log-log scale.
@@
Figure S8: Benchmarking molecule sampling strategies with GPT. Temperature, top-$k$, and top-$p$ sampling are used to generate molecules with the fine-tuned S4 models, in increasing parameters. Syntactic quality (a), structural similarity to the fine-tuning set (b), and internal diversity (c) are visualized. Same sampling and plotting parameters are used as Figure 5.

### v2

@@ Introduction
Temperature has the most substantial impact on the characteristics of de novo designs (Fig. 5). Higher temperature values increase the design diversity across datasets (Fig. 5a,c), but reduce validity and similarity to the fine-tuning set (Fig. 5a,b), in agreement with previous works[moret2023leveraging, ozccelik2024chemical]. For top-$k$ sampling, the smallest value of k ($k=3$) causes mode collapse, i.e., designs are valid but repetitive (Fig. 5a), suggesting that considering three candidate tokens might be insufficient to design diverse molecules. For $k\geq 5$, the sampling behavior approximates temperature sampling with $T=1$, indicating that the top-5 tokens already cover the most likely tokens. Top-$p$ sampling behaves similarly to top-k sampling. For $p\leq 0.9$, designs are valid but repetitive, while larger values of $p$ resemble temperature sampling ($T=1$). In the former case, few tokens cover the $p$ threshold, causing mode collapse; in the latter case, adding more tokens has little effect due to their low probability. We call this behavior, which is consistent across model architectures and training regimes (Fig. S9, S10, S11), ‘filtering trap’.
@@
These findings contrast with natural language generation: while top-$k$ and top-$p$ outperform temperature sampling to generate natural language, temperature sampling is the primary sampling strategy to control molecule diversity. To gain further insights, we generated 102,400 designs (using the fine-tuned LSTM for VDR). We analyzed the frequency of appearance of each token in the top-5 and top-90% of the distribution (Fig. 6a,b), along with their mean sampling probability (Fig. 6c). While probable tokens, such as carbon, are consistently likely across generation steps, most tokens rarely appear among the top-5 and top-90%. This skewed distribution differentiates molecule design from natural language generation, where tokens can be sampled among hundreds of thousands of candidates. In the molecular field, the number of tokens is inherently constrained due to the number of elements that can be used in drug-like molecules, leading to a higher concentration of likely candidates and a less diverse set of options at each generation step. Furthermore, training chemical language models enforces the validity of generations, e.g., every branch and ring opening has to be closed, further limiting the options at parts of the generation step. Overall, our analysis reveals a unique behavior of molecule generation, and exemplifies the gaps between natural language and drug discovery applications.
@@
Table 1: Summary of identified pitfalls, solutions, and guidelines. These considerations were divided based on the evaluation stage they pertain to.
Stage Pitfalls Solutions Recommendations
@@
Library similarity Library size. Metrics like FCD and FDD are dependent on library size, and decrease with increasing number of designs. Scaling up. Similarity metrics plateau for large libraries, making them suitable even with few reference bioactive molecules. Evaluating large libraries. Report FCD and FDD values for more than 100,000 designs.
@@
Molecule generation Token filtering. Considering a small token subset during molecule sampling (via top-k or top-p sampling) cause mode collapse (repetitive, low-diversity designs). Temperature sampling. Temperature sampling is the most effective strategy to control diversity and balance novelty vs. validity. Varying T values. Controlling diversity through token subsets is ineffective due to the unique behavior of molecule sampling. Temperature sampling should be used, with varying $T$ to tune lead optimization and diversity-focused exploration.
Conclusions
Robust evaluation pipelines are essential to identify and expand the boundaries of generative deep learning in drug discovery. Despite the topic of model benchmarking having garnered remarkable attention [brown2019guacamol, polykovskiy2020moses], to date, no standardized guidelines exist on what choices to make when evaluating generative models and their designs. Our large-scale analysis across targets, metrics, and generative deep learning approaches uncovered previously overlooked factors of the evaluation pipeline that can distort the outcome of generative deep learning projects (Table 1).
One key finding of this study is the confounding effect of the number of generated designs on model quality evaluation. This issue has significant implications, as it can lead to an over- or underestimation of relative model performance (and design quality). To mitigate this, metrics of similarity, internal diversity, and distribution-learning capabilities should always be compared across libraries of the same size, regardless of model setup or architecture. To ensure a robust assessment of generative models, we recommend reporting these metrics for libraries containing at least 105 designs. Additionally, whenever possible, analyzing how chosen metrics vary with library size could further highlight potential pitfalls in current evaluation practices.
This required library size is significantly larger than those used in most existing benchmarks and comparative studies, highlighting the need for a re-evaluation of model performance in light of these findings – particularly when comparing de novo designs from different studies. Moreover, generating and evaluating such large-scale de novo designs calls for more cost-efficient assessment strategies. In this work, we have identified computationally efficient measures of distributional similarity (FDD) and internal diversity (number of substructures) as one of those examples.
When exploring model-centric approaches for ranking de novo designs, model likelihood emerged as a strategy to balance exploration and exploitation. However, extreme likelihood values – either too high or too low – often correspond to redundant or low-quality designs. Our analyses further revealed that specific low-quality molecules tend to appear frequently, underscoring the need for effective filtering strategies. In this regard, combining model likelihood with design frequency provides a promising, model-informed ranking approach. Nevertheless, it remains unclear how model likelihood correlates with more complex molecular properties – such as bioactivity and toxicity – beyond simple measures of molecular similarity. To address this, we encourage the community to systematically report model likelihoods or their analogs for each selected design, helping to ‘illuminate the opaque box’ by clarifying what these likelihoods capture and how they can be more effectively leveraged.
This study also draws distinctions between generating ‘language of chemistry’ and natural language. Unlike natural language, where tokens can be chosen from a vast vocabulary, molecular generation is inherently constrained by the limited number of chemical elements and feasible substructures. As a result, model predictions tend to be more concentrated around a narrower set of high-likelihood candidates, leading to different challenges in ensuring diversity and exploration. In this context, an interesting direction is fragment-based molecular representations [noutahi2024safe, mastrolorito2025fragsmiles, cheng2023groupselfies], which increases the number of available tokens. This ’chemical word’ level representation (different from the atom-level representation that is routinely used for de novo design) might help strengthen the bridges between natural and chemical language processing. It remains to be determined whether this change would reflect in an increased benefit of different sampling strategies, an area that warrants further investigation.
Overall, we discovered ‘traps’ and solutions to evaluate generative drug discovery approaches. While we focused on fine-tuned chemical language models, our results are expected to be applicable to evaluate a variety of generative deep learning approaches, e.g., graph-based approaches or goal-directed design [popova2018deep, abate2023graph, brown2019guacamol], and stimulate further research on potential caveats on the evaluation of generative drug discovery approaches. Meanwhile, we expect this work to set new standards to evaluate and compare different generative approaches for drug discovery, as well as novel tools for practitioners to generate promising molecular libraries and effectively navigate them in search of novel bioactive matter.
@@ Methods
Pre-training set. 2,372,675 SMILES strings were obtained from ChEMBL v33 [gaulton2017chembl]. Salts were removed, and molecules composed only of C, H, O, N, S, P, F, Cl, Br, and I atoms were retained. The SMILES strings of the remaining compounds were sanitized, canonicalized, and the charge and stereochemistry annotations were removed. SMILES strings longer than 80 tokens were dropped. The final set consisting of 1,584,858 molecules was randomly divided into training ($n$ = 1,500,000), validation (nn= 40,000), and test splits ($n$ = 44,858).
Fine-tuning sets. The fine-tuning datasets were curated from ExCAPE-DB [sun2017excape]. The targets Dopamine Receptor D3 (DRD3), peptidyl-prolyl cis/trans isomerase (PIN1), and Vitamin D Receptor (VDR) were selected. The bioactive molecules available in the pre-training set were excluded, and the remaining molecules underwent the same pre-processing steps as described above. 320 training, 128 validation, and 128 test molecules were randomly sampled among the pre-processed strings. Random sampling was repeated five times with different random seeds, obtaining five fine-tuning sets per protein target.
@@
where $T$ is the temperature parameter and $y_{t}$ is the logit output by the model for the token $t$. Increasing the temperature value increases the uniformity of the distribution (uniform distribution for $T\rightarrow\infty$), while decreasing the temperature value decreases the randomness (Dirac distribution for $T\rightarrow 0$). When $T$ = 1, the so-called multinomial sampling (where the model output determines the probability of generating each token) is performed. We experimented with 0.5, 0.75, 1.0, 1.25, 1.5, 1.75, and 2.0 as $T$ values in this study.
Top-$k$ sampling. Top-$k$ sampling [holtzman2019curious] samples the next token from the most likely $k$ tokens. Increasing the $k$ values to the number of tokens in the vocabulary makes it equivalent to temperature sampling. Using $k=1$ is equivalent to greedy sampling, i.e., sampling the most likely token at each step. We experimented with values of $k$ equal to 3, 5, 10, 15, 20, 25, and 30 (with a vocabulary size equal to 33).
Top-$p$ sampling. Top-$p$ sampling [holtzman2019curious] (also known as nucleus sampling) samples the next token from the minimum cardinality set whose summed probabilities exceed the threshold $p$ ($0\leq p\leq 1$). Decreasing the value of $p$ includes fewer tokens in the selection and helps to avoid degenerate outputs, while trading diversity off [holtzman2019curious]. Top-$p$ sampling approximates greedy sampling as $p\rightarrow 0$ and is equivalent to temperature sampling when $p=1.0$. In this study, we experimented with values of $p$ equal to 0.5, 0.6, 0.7, 0.8, 0.9, 0.925, 0.950, and 0.975.
Sampling strategy. From each target (3x), each data split (5x), and each model architecture (3x), we generated 1,000,000 SMILES strings per sampling strategy and sampling parameter ($T$, $k$, and $p$, 22 values tested in total). This resulted in a total of $3\times 5\times 3\times 22\times 10^{6}=990,000,000=9.9\times 10^{8}\approx 10^{9}$ molecules that were designed and evaluated in this study.
@@
  • The Fréchet ChemNet Distance (FCD) [preuer2018frechet] was computed using the fcd library released by the authors.
  • The Fréchet distance [frechet1957distance] between molecular descriptor distributions (FDD) was computed using the following descriptors: octanol-wated partitioning coefficient [wildman1999prediction], molecular weight, number of hydrogen bond donors, number of rings, and topological surface area. Molecular descriptors were computed via rdkit, and min-max normalized to global maximum and minimum values before distance calculation.
  • The substructure similarity was computed via Tanimoto similarity on extended connectivity fingerprints (ECFPs) [rogers2010extended]. ECFPs were computed with rdkit (fpSize=2048, and radius=2).
Internal diversity The number of clusters was computed by using the sphere exclusion algorithm implemented in the LeaderPicker module of rdkit, which is equivalent to computing #Circles metric [xiemuch]. We used a distance threshold of 0.6 on the Tanimoto similarity on ECFPs. Number of substructures was calculated by counting the number of unique fingerprint keys identified by the Morgan algorithm (rdkit, radius=2).
@@
where $t$ runs over the tokens in the designed sequence. The log-sum-exp trick was used to mitigate numerical instabilities and log-likelihoods for each string were divided by the number of tokens.
Availability of data and materials
The Python code to replicate our study is on GitHub at the following URL: https://github.com/molML/jungle-of-generative-drug-discovery.
@@
Funding
This research was co-funded by the European Union (ERC, ReMINDER, 101077879). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The authors also acknowledge support from the Irene Curie Fellowship, the Centre for Living Technologies.
Author Contributions
Conceptualization: both authors. Data curation: R.Ö. Formal analysis: both authors. Investigation: both authors. Methodology: both authors. Software: R.Ö. Visualization: R.Ö. Writing – original draft: R.Ö. Writing – review and editing: both authors.
Acknowledgements
The authors thank Selen Parlar and Helena Brinkmann for their feedback on the manuscript.
@@ Acknowledgements
How Evaluation Choices Distort the Outcome of Generative Drug Discovery
@@
Figure S1: Structural similarity of held-out active and inactive molecules to fine-tuning sets. The structural similarity is quantified as Tanimoto similarity between Morgan fingerprints (radius=2, nBits=2048). The maximum similarity of each held-out molecule to the fine-tuning set is computed and visualized as a distribution.
Figure S2: Pretraining designs similarity and diversity.1,000,000 designs are generated via each pretrained model. FCD (a), FDD (b), uniqueness (c), number of clusters (d), and number of structures (e) are reported in increasing number of designs.
Figure S3: Diversity metrics divided by library size. Number of substructures and clusters are computed in increasing library sizes and divided by the number of molecules in the library. The same experimental and visualization setups are used as Figure 2.
Figure S4: Effect of synthesizability filtering on internal diversity. Synthetic accessibility of the designs is computed [ertl2009estimation]and the generations with a score above six are filtered out [ertl2009estimation]. Number of substructures is computed with the remaining compounds. The same experimental and visualization setups are used as Figure 2.
Figure S5: Speed comparison of internal diversity metrics. Number of clusters and substructures are computed in increasing library sizes, and the number of seconds per computation is measured. Each computation is repeated 10 times, and the average time passed is reported on a log-log scale.
@@
Figure S10: Benchmarking molecule sampling strategies with GPT. Temperature, top-$k$, and top-$p$ sampling are used to generate molecules with the fine-tuned GPT models, in increasing parameters. Syntactic quality (a), structural similarity to the fine-tuning set (b), and internal diversity (c) are visualized. Same sampling and plotting parameters are used as Figure 5.
Figure S11: Benchmarking molecule sampling strategies with pretrained LSTM. Temperature, top-$k$, and top-$p$ sampling are used to generate molecules with the pretrained models, in increasing parameters. Syntactic quality (a) and internal diversity (b) are visualized. Same sampling and plotting parameters are used as Figure 5.